The ratio of thermal conductivity of two rods different material is 5:4.The two rods of same area of cross section and same thermal resistance wil the length in the ratio:

Grade:12

1 Answers

askIITianexpert IITDelhi
14 years ago

Thermal resistance =l/KA

So, l1/l2 = K1/K2 =5/4